The vertex of the parabola below is at the point (2, 4) and the point (3, 6) is on the parabola. What is the equation of the parabola? A.y = 6(x - 2)2 + 4 B.y = 2(x - 2)2 + 4 C.x = 4(y - 3)2 + 1 D.y = 3(x - 4)2 + 2
Answer by stanbon(75887)

The vertex of the parabola below is at the point (2, 4) and the point (3, 6) is on the parabola. What is the equation of the parabola? A.y = 6(x - 2)2 + 4 B.y = 2(x - 2)2 + 4 C.x = 4(y - 3)2 + 1 D.y = 3(x - 4)2 + 2
----
A and B are candidates for being the correct solution
because f(2) = 4 for both.
---
B is the correct solution because f(3)=6
=====
Cheers,
Stan H.